OpenStudy (anonymous):

diana took her team to breakfast to celebrate sonias birthday.brians breakfast cost $9.50 steves breakfast cost $7.25.Both ken and judys breakfast cost $6.95 each.dianas breakfast cost $7.45.sonias breakfast being that it was her birthday was free.if the sales tax in palm beach county is 6% how much was the total bill including the sales tax??

OpenStudy (anonymous):

u want just the answer or how to do it

OpenStudy (anonymous):

or both

OpenStudy (jdoe0001):

\(\bf \begin{array}{llll} person&cost\\ \hline\\ brian&9.50\\ steve&7.25\\ ken&6.95\\ judy&6.95\\ sonia&0.00\\ \hline\\ &\square ? \end{array}\qquad 6\%\ of \ \square =\cfrac{6}{100}\times \square \\ \quad \\ \quad \\ \textit{total bill}=\square +\left(\cfrac{6}{100}\times \square \right) \)

OpenStudy (anonymous):

both

OpenStudy (anonymous):

u add the cost of every body and then multiply it by .06 which is $18.39

OpenStudy (anonymous):

add them 9.50 + 7.25 + 6.95 + 6.95 = bill without tax now take that total and multiply it by .06 (decimal form of 6%. That is your tax. Now add your tax to the total you got from adding earlier and that is your total bill.

OpenStudy (anonymous):

30.65 + .06 = 30.71 zrry my mistake

OpenStudy (anonymous):

isnt it 7.45 not 6.95

OpenStudy (anonymous):

it equal 31.15

OpenStudy (anonymous):

so the one that said 6.95 thats wrong is 7.45

OpenStudy (anonymous):

yes the answer is 37.15

OpenStudy (anonymous):

31.21 is the answer

OpenStudy (anonymous):

how u got 37.15

OpenStudy (anonymous):

add all of them and then add the answer to 6%

OpenStudy (anonymous):

add the answer to 0.06

OpenStudy (anonymous):

yeah 6% is the same thing

OpenStudy (anonymous):

no it not try it wit 6%

OpenStudy (anonymous):

it gives u 37.15

OpenStudy (anonymous):

9.50 + 7.25 + 6.95 + 6.95 + 7.45 = 38.10 (total without tax) 38.10 * .06 = 2.29 (tax) 38.10 + 2.29 = 40.39

OpenStudy (mathstudent55):

$9.50 $7.25 $6.95 $6.95 + $7.45 --------- $30.65 The total amount before tax is $38.10. This is 100% of the cost (before tax.) Since tax is 6% of the total, the bill will be 100% + 6% = 106% of the before-tax amount. That means you need to calculate 106% of $38.10. 106% of $38.10 = 1.06 * $38.10 = $40.386 = $40.39

OpenStudy (anonymous):

edolivera...you do not add to .06....you multiply by .06, then add the tax to the total without tax

OpenStudy (anonymous):

how did u get 30.65

OpenStudy (anonymous):

I think that was a mistake...because when you add them up, you get 38.10

OpenStudy (mathstudent55):

The amount of $30.65 is an error because that did not take into account the $7.45 of Diana's breakfast.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

final answer is $ 40.39

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

who work should i show on my paper

OpenStudy (anonymous):

which ever one you understand the best. You will arrive at the same answer either way.

OpenStudy (mathstudent55):

Use kellie's or mine. We both got the right answer using different approaches.

OpenStudy (anonymous):

ok but whih is easy for my teacher not to now

OpenStudy (anonymous):

which

OpenStudy (anonymous):

what do you mean by " which is easier for my teacher not to know " ?

OpenStudy (anonymous):

your teacher probably knows both ways

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

51.95 is it I think

OpenStudy (anonymous):

Hey kellegirl33 would it be 51.95

OpenStudy (anonymous):

no..it is 40.39

OpenStudy (anonymous):

Kk thanx

OpenStudy (anonymous):

anytime :)